# Extracting Affect Aggregates from Longitudinal Social Media Data with Temporal Adapters for Large Language Models

### v1

@@ Abstract
This paper proposes temporally aligned Large Language Models (LLMs) as a tool for longitudinal analysis of social media data. We fine-tune Temporal Adapters for Llama 3 8B on full timelines from a panel of British Twitter users, and extract longitudinal aggregates of emotions and attitudes with established questionnaires. We validate our estimates against representative British survey data and find strong positive, significant correlations for several collective emotions. The obtained estimates are robust across multiple training seeds and prompt formulations, and in line with collective emotions extracted using a traditional classification model trained on labeled data. To the best of our knowledge, this is the first work to extend the analysis of affect in LLMs to a longitudinal setting through Temporal Adapters. Our work enables new approaches towards the longitudinal analysis of social media data.
@@ Introduction
A number of recent studies has used Large Language Models (LLMs) to perform cross-sectional surveys in silico (e.g., Argyle et al. 2023; Bisbee et al. 2023; Durmus et al. 2023), modeling human-based survey responses by utilizing LLMs. While showing great promise, such prompt-based in silico surveys have not been aligned well from a temporal perspective, meaning that LLM training data and survey fieldwork periods typically did not meaningfully overlap in terms of time. Independently, many studies have investigated affect aggregates from social media data, in particular collective emotions (e.g., Golder and Macy 2011; Pellert et al. 2022; Metzler et al. 2023). However, these methods typically rely on expensively curated dictionaries or on labeled data.
@@
In this paper, we present a novel method for extracting longitudinal affect aggregates by using Temporal Adapters fine-tuned on user-generated data from social media, as illustrated in Figure 1. We base our analysis on a panel of 24,000 British Twitter users and obtain their full timelines from November 2019 to June 2020. First, we fine-tune Temporal Adapters for Llama 3 8B on 7-day subsets of the Twitter data. Second, we prompt the fine-tuned model with multiple established questionnaires (YouGov 2024a, c) and extract longitudinal affect aggregates from token probabilities. We evaluate our results against survey data of the British adult population and find strong positive and significant correlations for a subset of collective emotions. A comparison of our method with a traditional classification model trained on labeled data produces strongly agreeing results. We demonstrate robustness against prompt variation with a survey instrument that measures the same affective phenomena (Clark and Watson 1994). Experiments with synthetically mixed data indicate the internal validity of our method. We exemplary show that our method can be used for the extraction of more complex collective attitudes as well.
@@
The main contribution of this work is that it extends previous in silico surveys with LLMs to a longitudinal setting in which survey data and LLM training data are temporally aligned. Our method is less sensitive to strata-dependent biases embedded in pre-trained LLMs since we fine-tune LLMs on user-generated data and focus on longitudinal changes in affect aggregates. Compared to previous methods that extract affect aggregates from user-generated data, our approach is not fixed to a specific survey question and neither relies on expensively curated dictionaries, nor on labeled data. We provide a Python implementation under MIT license alongside our paper to facilitate replication and future work on extracting affect aggregates with LLMs [^1] https://github.com/dess-mannheim/temporal-adapters.
Related Work
Attitudes, Opinions, and Values in LLMs
@@
Many recent studies have investigated the application of longitudinal surveys to LLMs (e.g. Argyle et al. 2023; Bisbee et al. 2023; Atari et al. 2023; Von Der Heyde, Haensch, and Wenz 2023) Researchers typically prompted LLMs with survey questions from existing questionnaires that were originally developed to survey human populations (Ma et al. 2024; Agnew et al. 2024). This setup assumes that the diverse data on which LLMs were trained enables them to be viable proxies for population-level estimates of individual attributes. However, most research neglects the temporal alignment of training data and survey data, i.e., the fieldwork periods in which survey data was collected. We propose Temporal Adapters fine-tuned on user-generated data with a known date of creation to improve longitudinal alignment.
@@
Research has shown that openly available, pretrained LLMs produce estimates that can be both culturally and politically biased when compared to population-level survey data (e.g., Motoki, Pinho Neto, and Rodrigues 2023; Santurkar et al. 2023; Hartmann, Schwenzow, and Witte 2023; Adilazuarda et al. 2024). Since our method is based on fine-tuning an existing LLM, we likely inherit some of these biases. Still, by fine-tuning the model on user-generated data, we control the sampling process of whose attitudes, opinions, and values the model will learn.
@@
LLMs were found to be sensitive to survey wording (Tjuatja et al. 2024; Dominguez-Olmedo, Hardt, and Mendler-Dünner 2023; Röttger et al. 2024; McIlroy-Young et al. 2024), and the particular answer scoring method (Wang et al. 2024b, a). We prompt each answer option separately to circumvent order effects and we use an answer-prefix to be less dependent on first-token probabilities. We demonstrate robustness on prompt wording by extracting collective emotions with multiple survey instruments.
@@
LLMs used for surveys are so far either studied as they were pre-trained, or they are fine-tuned on survey data (Ramezani and Xu 2023; Kim and Lee 2023). With pre-trained models, researchers typically use silicon sampling and prepend a diverse set of persona description, often derived from existing survey data, to the survey question (Argyle et al. 2023; Sun et al. 2024). A major drawback of silicon sampling is that is purely relies on a model’s pre-training data, potentially mis-portraying parts of society (Wang, Morgenstern, and Dickerson 2024), and with no way of learning from the attitudes, opinions, and values that they voice. Our method is instead based around incorporating data generated by specific users at a specific time into LLMs used for surveys.
Kim and Lee (2023) added embeddings obtained from LLMs to a neural network to predict longitudinal changes in opinions. We instead work directly with a state-of-the-art LLM, to leverage its general language capabilities, and to extract affect aggregates using the same survey question that was asked to human survey participants. Ramezani and Xu (2023) trained an LLM on existing survey data and were able to improve similarity to human survey responses. The training of Temporal Adapters is independent of existing survey data and allows us to extract affect aggregates from specific user generated datasets.
Measuring Affect in Surveys
Our primary focus in this paper is on extracting emotion aggregates from user-generated social media data, even if our setup does not prescribe which survey question we use to extract collective affect, i.e., which construct we want to study. In social psychology and sociology, affect is used as the most general term that encompasses phenomena such as emotions and moods (Mohiyeddini and Bauer 2013; Rogers and Robinson 2014). Emotions are typically more specific and short-lived, lasting between minutes and days, while moods describe phenomena that are more diffuse and last between hours and weeks (Oatley and Johnson-Laird 2014), but the temporal distinction is not completely clear-cut.
@@
Researchers generally distinguish between felt and expressed emotions or moods (Rogers and Robinson 2014). Neither surveys nor social media data can tap into internal, felt emotions directly, so we focus only on expressed emotions in this paper. Survey measures of affect are further influenced by recall bias and by reactivity because of their retrospective nature. Reactivity refers to the fact that ‘the very knowledge that one is being observed can alter emotional experience’ (Rogers and Robinson 2014, p. 286). Survey participants might also be more or less willing to report different emotions. Affect is an inherently social phenomena and central to interpersonal communication (Lively and Weed 2016), so social media platforms might intuitively seem like the perfect source of found data on expressed emotions. In contrast to surveys, posting on social media can be more immediate and less distanced from the affective experience itself. Working with found data is also considerably cheaper than surveying.
Social Media Affect Macroscopes
Text data generated by users on social media has been used to study election outcomes (Gayo-Avello 2013) and consumer confidence (Pasek et al. 2018), with challenging results. One of the most common applications is the extraction of emotion macroscopes (Golder and Macy 2011; Garcia et al. 2021; Pellert et al. 2022), in particular in response to catastrophic events (Garcia and Rimé 2019; Jones and Silver 2020) such as the beginning of the COVID-19 pandemic in 2020 (e.g., Valdez et al. 2020; Lwin et al. 2020; Ashokkumar and Pennebaker 2021; Metzler et al. 2023).
Similar to our setup, estimates are aggregated for specific groups, or on a population level. In contrast to previous approaches, our method does not rely on labeled training data or on expensively created dictionaries.
@@
Social media data suffers from sampling biases and performative behavior of users due to platform effects and community norms (Sen et al. 2021). Estimates obtained from social media samples often do not generalize to a target population because of differences in internet penetration and platform use (Amaya et al. 2020). In addition, behavioral differences between groups lead to biases in representation (Olteanu et al. 2019). In other words, different groups are more or less inclined to express emotions or attitudes on social media. Still, previous studies found strongly positive and robust correlations between collective emotions extracted from social media and survey responses on a population level (Garcia et al. 2021; Pellert et al. 2022). We extend this research by introducing an extraction method that is more flexible by extracting constructs directly through prompting LLMs with survey questions.
@@
Longitudinal Datasets
We extract affect aggregates from Twitter panel data for Great Britain, covering 35 weeks from November 2019 to June 2020. This time frame includes both New Year 2020 as well as the first UK COVID-19 lockdown on March 23rd, 2020. It allows us to investigate both seasonal patterns and a catastrophic event that had a large impact on emotions and attitudes, as identified in previous research. We decided against a larger time frame due to the large amount of computational resources required to fine-tune our model.
Twitter Panel Data
We create a panel of in total 21,576 Twitter users (13.6 million tweets) with the following method: We use the commercial service Brandwatch to sample 10,000 tweets per day from accounts in Great Britain between the beginning of 2019 and March 2023. Brandwatch determines users’ locations based on profile information and geo-tagged data (Brandwatch 2020b). Next, we identify individual users in that sample and removed accounts that were classified as organizations as well as accounts with too low or too high activity levels. We select 24,000 users at random, half of which Brandwatch classified as ‘female’ and half of which were classified as ‘male’ based on self-reported profile information (Brandwatch 2020a). We retrieve their full timelines of tweets including retweets and replies back to the beginning of 2019. On average, for the 35 weeks from November 2019 to June 2020, we have 385,000 tweets per week available.
Questionnaires and Survey Data
We extract affect aggregates by querying an LLM with existing survey questionnaires and we compare our results to publicly available survey data that was collected using the same questions. The survey data was collected from a British online panel with a target population of all British adults aged 18+ (YouGov 2024b). The panel provider uses active sampling and post-survey adjustment weights on age, gender, social class, region, and level of education.
@@
This question uses a multi-code answer scale, i.e., multiple of the following answer options can be selected: ‘happy’, ‘sad’, ‘energetic’, ‘apathetic’, ‘inspired’, ‘frustrated’, ‘optimistic’, ‘stressed’, ‘content’, ‘bored’, ‘lonely’, and ‘scared’. We exclude the instructions on how to answer as well as the answer options ‘other’ and ‘don’t know’ from our setup, as these are tailored to human survey response behavior. We assess our results by comparing them with aggregate survey data gathered from the above described British panel on the same question. Survey data is available in weekly waves as aggregates over 1890 to 2081 participants per wave.
@@
To what extend have you felt [adjective] during the past week?
We focus on the two emotions included in both the PANAS-X and in YouGov’s survey data, scared and sad. The answer options are ‘very slightly or not at all’, ‘a little’, ‘moderately’, ‘quite a bit’, and ‘extremely’, to be answered for each adjective – in contrast to YouGov’s answer options, which are directly multiple choice among emotions. We extract answer probabilities for each adjective and each answer option, and combine all of them into a single score for each emotion, according to the instructions provided by Clark and Watson (1994). We compare our results to survey data gathered by YouGov in Britain’s Mood, Measured Weekly for the respective emotions.
The National Health Service (NHS)
In addition to collective emotions, we also extract a collective attitude towards the NHS with the following question (YouGov 2024c):
@@
We spell out the abbreviation ‘NHS’ since we query for this survey question without additional context. We include the answer options ‘get better’ and ‘get worse’ into our analysis and compare our results to survey data from YouGov on the same question. Survey data is available in monthly waves as aggregates over 1618 to 1817 participants per wave.
Figure 2: Temporal Adapter Fine-Tuning. We concatenate each week’s tweets into chunks for batch-based fine-tuning. While fine-tuning each week’s parameter-efficient LoRA (Hu et al. 2021) adapter, the original model weights are kept frozen. Fine-Tuning is performed with the causal language modeling task, i.e., next-token prediction.
Temporal Adapters for LLMs
Figure 1 provides an overview of our proposed setup, comprising two separate steps: First, we fine-tune Temporal Adapters for Llama 3 8B on the Twitter panel data described in the previous section. Second, we prompt the model with one of the selected survey questions and one weekly adapter activated at a time. We extract token probabilities for each answer option across all weeks, which we combine into longitudinal macroscopes of mood and attitudes.

### v2

@@ Abstract
This paper proposes temporally aligned Large Language Models (LLMs) as a tool for longitudinal analysis of social media data. We fine-tune Temporal Adapters for Llama 3 8B on full timelines from a panel of British Twitter users, and extract longitudinal aggregates of emotions and attitudes with established questionnaires. We focus our analysis on the beginning of the COVID-19 pandemic that had a strong impact on public opinion and collective emotions. We validate our estimates against representative British survey data and find strong positive, significant correlations for several collective emotions. The obtained estimates are robust across multiple training seeds and prompt formulations, and in line with collective emotions extracted using a traditional classification model trained on labeled data. We demonstrate the flexibility of our method on questions of public opinion for which no pre-trained classifier is available. Our work extends the analysis of affect in LLMs to a longitudinal setting through Temporal Adapters. It enables flexible, new approaches towards the longitudinal analysis of social media data.
@@ 1 Introduction
Several recent studies have used Large Language Models (LLMs) to perform cross-sectional surveys in silico (e.g., Argyle et al. 2023; Bisbee et al. 2023; Durmus et al. 2023), modeling human-based survey responses synthetically. While showing great promise, such prompt-based in silico surveys have not been aligned well from a temporal perspective, meaning that LLM training data and survey fieldwork periods (waves) typically did not meaningfully overlap temporally. Independently, many studies have investigated affect aggregates from social media data, in particular collective emotions (e.g., Golder and Macy 2011; Pellert et al. 2022; Metzler et al. 2023). Research on collective emotions is important as it can provide actionable insights for public health messaging in times of crisis (Lwin et al. 2020) and on the social resilience of communities (Garcia and Rimé 2019). However, these methods don’t scale well as they typically rely on expensively curated dictionaries or on labeled training data.
@@
In this paper, we present a novel method for extracting longitudinal affect aggregates by using Temporal Adapters fine-tuned on user-generated data from social media, as illustrated in Figure 1. Adapters allow researchers to efficiently fine-tune LLMs with data from new contexts by training only a very small fraction of additional parameters while keeping the original model weights frozen. We base our analysis on a panel of 21,576 British Twitter users and obtain their full timelines from November 2019 to June 2020. First, we fine-tune Temporal Adapters for Llama 3 8B on 7-day subsets of the Twitter data. Second, we prompt the fine-tuned model with multiple established questionnaires (YouGov 2024a, e) and extract longitudinal affect aggregates from token probabilities. We evaluate our results against survey data collected from a representative sample of the British adult population and find strong positive and significant correlations for a subset of collective emotions. A comparison of results obtained using our method with a traditional classification model trained on labeled data shows strong agreement. We demonstrate robustness against prompt variation with an additional survey instrument that is constructed to measure the same affective phenomena (Clark and Watson 1994). Experiments with synthetically mixed data indicate the internal validity of our method. Finally, we demonstrate that our flexible method can be used for the extraction of more complex collective attitudes as well.
@@
The main contribution of this work is that it extends previous approaches to in silico surveys with LLMs to a longitudinal setting in which survey data and LLM training data are temporally aligned. Compared to previous methods that extract affect aggregates from user-generated data, our approach is not fixed to a specific survey question and neither relies on expensively curated dictionaries, nor on labeled data. Instead, our method allows researchers make use of the wealth of established survey questionnaires to extract affect aggregates from social media data. We provide a Python implementation under MIT license alongside our paper to facilitate replication and future work on extracting affect aggregates with LLMs. [^1] https://github.com/dess-mannheim/temporal-adapters
2 Related Work
2.1 Attitudes, Opinions, and Values in LLMs
@@
Many recent studies have investigated the application of cross-sectional surveys to LLMs (e.g. Argyle et al. 2023; Bisbee et al. 2023; Atari et al. 2023; Von Der Heyde, Haensch, and Wenz 2023). Researchers typically prompt LLMs with survey questions from existing questionnaires that were originally developed to survey human populations (Ma et al. 2024; Agnew et al. 2024). This setup assumes that the diverse data on which LLMs were trained enables them to be viable proxies for population-level estimates of individual attributes. However, most research neglects the temporal alignment of training data and survey data, i.e., the fieldwork periods in which survey data was collected. We propose Temporal Adapters fine-tuned on user-generated data with a known date of creation to enable longitudinal alignment.
@@
Research has shown that openly available, pretrained LLMs produce estimates that can be both culturally and politically biased when compared to population-level survey data (e.g., Motoki, Pinho Neto, and Rodrigues 2023; Santurkar et al. 2023; Hartmann, Schwenzow, and Witte 2023; Adilazuarda et al. 2024). Since our method is based on fine-tuning an existing LLM, we likely inherit some of these biases. Still, by fine-tuning the model on user-generated data, we have some degree of control the sampling process of whose attitudes, opinions, and values the model will learn.
@@
LLMs were found to be sensitive to survey wording (Tjuatja et al. 2024; Dominguez-Olmedo, Hardt, and Mendler-Dünner 2023; Röttger et al. 2024; McIlroy-Young et al. 2024), and the particular answer scoring method (Wang et al. 2024b, a). We prompt each answer option separately to counteract order effects and we use an answer-prefix to be less dependent on first-token probabilities. We demonstrate that our approach is robust with respect to specific prompt wording by extracting collective emotions with multiple survey instruments.
Temporal Adaptation of Language Models
Besides surveys, temporal adaptation of language models has been studied to combat language change over time. Röttger and Pierrehumbert (2021) used social media data to temporally adapt language models to both unlabeled and labeled training data. They found that, for document classification, temporal adaptation does not provide additional benefits to domain adaptation. Similarly, Lazaridou et al. (2021) create a dynamic benchmark to account for changes in language when evaluating the performance of language models, and Amba Hombaiah et al. (2021) propose incremental training regimes for language models. Instead of being concerned with general language model performance over time, our work focuses on using Temporal Adapters to extract affect aggregates from longitudinal social media data.
@@
LLMs used for surveys are so far either studied as they were pre-trained, or they are fine-tuned on survey data (Ramezani and Xu 2023; Kim and Lee 2023). With pre-trained models, researchers typically use silicon sampling and prepend a diverse set of persona description, often derived from existing survey data, to the survey question (Argyle et al. 2023; Sun et al. 2024). A major drawback of silicon sampling is that it is purely based on a model’s pre-training data, potentially mis-portraying parts of society (Wang, Morgenstern, and Dickerson 2024). Our method is instead based around incorporating data generated by specific users at a specific time into LLMs that are then used for surveying.
Kim and Lee (2023) added embeddings obtained from LLMs to a neural network to predict longitudinal changes in opinions. We instead work directly with a state-of-the-art LLM, to leverage its general language capabilities, and to extract affect aggregates using the same survey question that was asked to human survey participants. Ramezani and Xu (2023) trained an LLM on existing survey data and were able to improve similarity to human survey responses. The training of Temporal Adapters is independent of existing survey data and allows us to extract affect aggregates from specific user-generated text datasets.
2.2 Measuring Affect in Surveys
With our approach, we extract affect aggregates from user-generated social media data. Our setup does not prescribe which exact survey question we use, flexibly allowing us to choose which construct specifically we want to study. In social psychology and sociology, affect is used as the most general term that encompasses phenomena such as emotions and moods (Mohiyeddini and Bauer 2013; Rogers and Robinson 2014). Emotions are typically more specific and short-lived, lasting between minutes and days, while moods describe phenomena that are more diffuse and last between hours and weeks (Oatley and Johnson-Laird 2014), but the temporal distinction is not completely clear-cut.
@@ 2.2 Measuring Affect in Surveys
Researchers generally distinguish between felt and expressed emotions or moods (Rogers and Robinson 2014). Neither surveys nor social media data can tap into internal, felt emotions directly, so we focus only on expressed emotions in this paper. Survey measures of affect are further influenced by recall bias and by reactivity because of their retrospective nature. Reactivity refers to the fact that ‘the very knowledge that one is being observed can alter emotional experience’ (Rogers and Robinson 2014, p. 286). Survey participants might also be more or less willing to report different emotions. Affect is an inherently social phenomena and central to interpersonal communication (Lively and Weed 2016), so social media platforms might intuitively seem like a suitable source of found data containing expressed emotions. In contrast to surveys, posting on social media can be more immediate and less distanced from the affective experience itself. Working with found data is also considerably cheaper than surveying. Disadvantages include that people on social media may behave vis-a-vis an audience, creating a source of performative bias.
2.3 Social Media Affect Macroscopes
Text data generated by users on social media has been used to study election outcomes (Gayo-Avello 2013) and consumer confidence (Pasek et al. 2018), with challenging results. One of the most common applications is the extraction of emotion macroscopes (Golder and Macy 2011; Garcia et al. 2021; Pellert et al. 2022), in particular in response to catastrophic events (Garcia and Rimé 2019; Jones and Silver 2020) such as the beginning of the COVID-19 pandemic in 2020 (e.g., Valdez et al. 2020; Lwin et al. 2020; Ashokkumar and Pennebaker 2021; Metzler et al. 2023). Collectively experienced emotions can be of different quality, magnitude, and duration than individual emotions (Goldenberg et al. 2020) and research on collective emotions can provide actionable insights for public health messaging in times of crisis (Lwin et al. 2020) and on the social resilience of communities (Garcia and Rimé 2019). These methods are based on expensively created dictionaries or on models trained with labeled data (e.g., Mohammad et al. 2018), which are then typically used to detect positive/negative sentiment or individual emotions. Similar to our setup, estimates are aggregated for specific groups, or on a population level. In contrast to such earlier approaches, our method scales well by not relying on labeled training data or on expensively created dictionaries and can be flexibly applied to any survey question.
@@ 2.3 Social Media Affect Macroscopes
As mentioned before, social media data can suffer from sampling biases and consequences of performative behavior of users due to platform effects and community norms (Sen et al. 2021). Estimates obtained from social media samples often do not generalize to a target population because of differences in internet penetration and platform use (Amaya et al. 2020). In addition, behavioral differences between groups lead to biases in representation (Olteanu et al. 2019). In other words, different groups are more or less inclined to express emotions or attitudes on social media. Still, previous studies found strongly positive and robust correlations between collective emotions extracted from social media and survey responses on a population level (Garcia et al. 2021; Pellert et al. 2022). We extend this research by introducing an extraction method that is more flexible by extracting constructs directly through prompting LLMs with survey questions.
@@
3 Longitudinal Datasets
We extract affect aggregates from Twitter panel data for Great Britain, covering 35 weeks from November 2019 to June 2020. This time frame includes both New Year 2020 and the first UK COVID-19 lockdown on March 23rd, 2020. It allows us to investigate both seasonal patterns and a catastrophic event that had a large impact on emotions and attitudes, as identified in previous research. We decided against a larger time frame because of the large amount of computational resources required to fine-tune our model. The selected time frame follows previous research on extracting collective emotions from social media data during the COVID-19 pandemic (e.g., Metzler et al. 2023), but we would also like to emphasize the uniqueness of the situation. Substantive results such as cross-correlations between individual emotions might not generalize to other points in time, but drastically changing emotions offer a unique opportunity for our methodological research.
3.1 Twitter Panel Data
We create a panel of in total 21,576 Twitter users (13.6 million tweets) with the following method: We use the commercial service Brandwatch to sample 10,000 tweets per day from accounts in Great Britain between the beginning of 2019 and March 2023. Brandwatch determines the location of users based on profile information and geo-tagged data (Brandwatch 2020b). Next, we identify individual users in that sample and remove accounts that were classified as organizations, as well as accounts with too low or too high activity levels. We select the same number of users classified by Brandwatch as ‘female’ and ‘male’, based on self-reported profile information (Brandwatch 2020a). We retrieve their full timelines of tweets including retweets and replies back to the beginning of 2019. On average, for the 35 weeks from November 2019 to June 2020, we have 385,000 tweets per week available.
3.2 Questionnaires and Survey Data
We extract affect aggregates by querying an LLM with existing survey questionnaires and compare our results to publicly available survey data that was collected using the same questions. The survey data was collected from a British online panel with a target population of all British adults aged 18+ (YouGov 2024d). The panel provider uses active sampling and post-survey adjustment weights based on age, gender, social class, region, and level of education.
@@ 3.2 Questionnaires and Survey Data
This question uses a multicode answer scale, that is, multiple of the following answer options can be selected: ‘happy’, ‘sad’, ‘energetic’, ‘apathetic’, ‘inspired’, ‘frustrated’, ‘optimistic’, ‘stressed’, ‘content’, ‘bored’, ‘lonely’, and ‘scared’. We assess our results by comparing them with aggregate survey data gathered from the aforementioned British panel on the same question. Survey data is available in weekly waves as aggregates over 1890 to 2081 participants per wave.
@@
To what extent have you felt [adjective] during the past week?
We focus on two emotions included in both the PANAS-X and in YouGov’s survey data, scared and sad. The answer options are ‘very slightly or not at all’, ‘a little’, ‘moderately’, ‘quite a bit’, and ‘extremely’, to be answered for each adjective – in contrast to YouGov’s answer options, which are directly multiple choice among emotions. We extract answer probabilities for each adjective and each answer option, and combine all of them into a single score for each emotion, according to the instructions provided by Clark and Watson (1994). We compare our results to survey data gathered by YouGov in Britain’s Mood, Measured Weekly for the respective emotions.
Collective Public Opinion Beyond Emotions
In addition to collective emotions, we include 3 public opinion questions into our analysis to demonstrate the flexibility of our approach. We select these questions based on available YouGov data for evaluation and based on their relevance to the beginning of the COVID-19 pandemic, during which the UK government had to implement stringent measures with large consequences on people’s everyday life. We extract collective attitudes towards the National Health Service (NHS) with the following question (YouGov 2024e):
@@
We spell out the abbreviation ‘NHS’ since we query for this survey question without additional context. We further extract public attitudes towards Boris Johnson as a Prime Minister with the following question (YouGov 2024c):
Do you think that Boris Johnson is doing well or badly as Prime Minister?
And we evaluate public approval of the UK government with the following question (YouGov 2024b):
Do you approve or disapprove of the Government’s record to date?
We include all respective answer options into our analysis and compare our results to survey data by YouGov for the respective question. Survey data is available in monthly waves for attitudes towards the NHS and towards Boris Johnson, and in weekly waves for Government Approval, with varying survey dates and varying numbers of participants.
4 Temporal Adapters for LLMs
Figure 1 provides an overview of our proposed setup, comprising two separate steps: First, we fine-tune Temporal Adapters for Llama 3 8B on the Twitter panel data described in the previous section. Second, we prompt the model with one of the selected survey questions and one weekly adapter activated at a time. We extract token probabilities for each answer option across all weeks, which we combine into longitudinal macroscopes of emotions and attitudes.